# Knowledge Homophily in Large Language Models

### v1

@@ 4.3. Homophily-guided Knowledge Retrieval
The reader model GPT-3.5 generates answers strictly based on the retrieved triples, and their correctness is evaluated by GPT-4. In Table 2, M/G-BS outperforms Baseline across all datasets and hops. Crucially, GNN-based beam search (G-BS) surpasses its homophily-agnostic MLP-based counterpart (M-BS), confirming the advantage of structural knowledge homophily in estimating knowledgeability. On 2-hop, G-BS has an improvement of 4.57% over the baseline, with larger gains on general KGs (e.g., T-Rex) than domain-specific ones (e.g., PharmKG). For 3-hop questions, all methods drop due to longer chains and semantic drift, but G-BS still has an improvement of 2.62% over the baseline. This superior performance of G-BS holds across different budgets for training the knowledgeability estimator in Figure 6 in Appendix.
@@ 5. Conclusion
Motivated by the structural knowledge organization of the human brain, this work explores the homophily structure of neural knowledge stored in LLMs, showing that LLMs’ knowledge about entities is strongly correlated with their neighbors in a knowledge graph. Building on this insight, we design a GNN regressor to estimate entity-level knowledgeability scores by leveraging their local neighborhood scores. We demonstrate the utility of these scores in two applications: selecting less-known triplets for efficient knowledge injection through fine-tuning, and improving retrieved context for enhancing multi-hop question answering.
@@ 6. Ethical Considerations
The knowledge homophily, where topologically proximate entities tend to share similar knowledgeability, increases the risk of knowledge-extraction attacks. An adversary can exploit homophily by crafting queries that target cohorts of related entities (or by combining entities across neighborhoods) to maximize coverage and force the model to reveal more information than intended. This raises privacy and copyright concerns because grouping semantically related entities makes it easier for the model to reconstruct or expose sensitive facts. To mitigate this risk, we propose defensive measures such as pre-filtering and sanitizing queries to remove suspicious or overly targeted prompts, rate-limiting repetitive/cohort queries, applying prompt-level heuristics that block requests for verbatim proprietary content, and deploying detector models or red-teaming workflows to identify adversarial extraction patterns. Combining these defenses with selective fine-grained access controls and differential privacy constraints can substantially reduce the attack surface introduced by knowledge homophily.
@@ A.2.3. Sensitivity Analysis: Knowledge Retrieval
To evaluate the robustness of our knowledge-aware beam search method (G-BS and M-BS), we conducted a sensitivity analysis on the amount of training data used for the knowledgeability estimators. The primary experiment in our paper utilizes GNN and MLP models trained on 40% of the available entities. This analysis investigates how performance on the multi-hop question-answering task varies when this training budget is reduced. For this sensitivity analysis we select the CoDEx-S dataset. We trained a series of GNN and MLP knowledgeability estimators on progressively larger subsets of entity data: 1%, 2%, 5%, 10%, 20%, and 40%. Each resulting estimator was then integrated into the G-BS and M-BS retrieval methods, respectively, and evaluated on the fixed set of 1,000 2-hop and 3-hop questions. The performance of the Semantic Beam Search baseline is independent of this training budget and remains constant. The results are presented in the Figure 6 for 2-hop (left) and 3-hop (right) QA performance. Our key findings of this analysis are:

### v2

@@ 4.3. Homophily-guided Knowledge Retrieval
Using a GPT-3.5 reader (restricted to retrieved triples) evaluated by GPT-4, we find that M/G-BS consistently outperforms the Baseline, as shown in Table 2. Crucially, G-BS surpasses the homophily-agnostic M-BS, validating the advantage of structural knowledge homophily. G-BS achieves a 4.57% improvement on 2-hop queries (favoring general KGs). While performance declines for all methods on 3-hop queries due to semantic drift, G-BS still secures a 2.62% improvement. These gains remain consistent across training budgets, as shown in Figure 6 in Appendix.
@@ 5. Conclusion
Inspired by the structural knowledge organization in the human brain, we investigate homophily in LLMs’ neural knowledge and validate it via correlated knowledgeability scores among neighboring entities in knowledge graphs. Building on this observation, we propose a GNN-based regressor that exploits local neighborhoods to estimate entity-level knowledgeability. We verify its effectiveness in selecting less-known triplets for efficient knowledge injection via fine-tuning, and in improving retrieval for multi-hop question answering. In the future, we plan to explore uncertainty-aware knowledge verification and dynamic homophily modeling to capture how homophily evolves as LLMs acquire new information.
Acknowledgements
This research is supported by the National Science Foundation (NSF) under grant number IIS 2524379 and NAIRR 250188.
@@ 6. Ethical Considerations
Knowledge homophily, where topologically proximate entities exhibit similar knowledgeability, amplifies the risk of knowledge-extraction attacks. Adversaries can exploit this structure by crafting queries over cohorts of related entities, thereby maximizing unintended information disclosure. This poses privacy and copyright risks, as semantically clustered entities facilitate reconstruction of sensitive facts. To mitigate these threats, several defensive strategies can be employed, including query pre-filtering and sanitization, rate-limiting cohort-based requests, prompt-level heuristics that block verbatim proprietary content, and detector or red-teaming mechanisms for identifying adversarial extraction patterns. When combined with fine-grained access control and differential privacy constraints, these measures can substantially reduce the attack surface introduced by knowledge homophily.
@@ A.2.2. Sensitivity Analysis: Knowledge Injection
Table 4. Statistics of the original knowledge graph and the sampled largest connected component.
Dataset # Nodes # Triplets # Avg. Deg # Avg. CC
Original Sampled Original Sampled Original Sampled Original Sampled
T-Rex 3153568 46891 6566790 193781 4.16 8.26 0.1473 0.5170
WD50K 41334 5140 233838 34208 11.31 13.31 0.0996 0.1332
PharmKG8K 7262 6877 479902 98537 132.16 28.65 0.2512 0.0824
MVPKG 137117 9055 1857410 255697 12.46 28.24 0.0013 0.0140
MVPKG w/o t 137117 9055 1857410 116127 12.46 12.82 0.0013 0.0140
CoDEx-S 2034 36543 35.93 0.0952
@@ A.2.3. Sensitivity Analysis: Knowledge Retrieval
To evaluate the robustness of our knowledge-aware beam search method (G-BS and M-BS), we conducted a sensitivity analysis on the amount of training data used for the knowledgeability estimators. The primary experiment in our paper utilizes GNN and MLP models trained on 40% of the available entities. This analysis investigates how performance on the multi-hop question-answering task varies when this training budget is reduced. For this sensitivity analysis, we select the CoDEx-S dataset. We trained a series of GNN and MLP knowledgeability estimators on progressively larger subsets of entity data: 1%, 2%, 5%, 10%, 20%, and 40%. Each resulting estimator was then integrated into the G-BS and M-BS retrieval methods, respectively, and evaluated on the fixed set of 1,000 2-hop and 3-hop questions. The performance of the Semantic Beam Search baseline is independent of this training budget and remains constant. The results are presented in the Figure 6 for 2-hop (left) and 3-hop (right) QA performance. Our key findings of this analysis are:
@@
A.2.4. Role of Knowledge Graph Probing
Our goal in this work is not to propose a new knowledge-graph (KG) probing method, but to use KG structure as a lens to study how an LLM’s factual knowledge is organized topologically, specifically, whether topologically proximate entities exhibit similar knowledgeability (“knowledge homophily”). Because the main scientific claim concerns the presence and utility of such a homophily pattern, the probing procedure is used solely as a measurement instrument to obtain a consistent entity-level knowledgeability signal that can be placed on the KG for analysis. As a result, the specific choice of probing method is not the focus of this work, and we do not claim superiority or equivalence across probing techniques. Our conclusions are based on the observed topological patterns induced by the chosen signal, rather than on properties of the probing method itself.
A.2.5. Scalability and Computational Considerations
While our experiments focus on small to mid-scale knowledge graphs, the computational implications for larger graphs warrant clarification. As we exploit the property of knowledge homophily: topologically close entities tend to exhibit similar knowledgeability, our approach does not require probing all entities or triplets in the graph, unlike prior methods that rely on exhaustive or iterative probing to identify knowledge deficiencies in LLMs (Song et al., 2025). Instead, we probe only a subset of entities, using their observed knowledgeability signals to infer the remaining ones via graph structure. This substantially reduces computational cost, as well as LLM-specific overhead such as inference time and token usage, since fewer prompts are required.
We acknowledge that our method introduces additional components beyond probing, namely the use of a GNN for knowledgeability inference over the graph. However, GNN inference is known to scale linearly with respect to the number of nodes and edges in the graph for a fixed number of layers, which is well established in the literature (Kipf, 2016; Veličković et al., 2017). Moreover, existing large-scale benchmarks have demonstrated that GNNs can be trained and evaluated on graphs with millions of entities, such as those in the Open Graph Benchmark (Hu et al., 2020). In comparison, the knowledge graphs used in our work are considerably smaller and therefore well within the regime where GNN-based inference is computationally feasible.
A.2.6. Datasets
Our experiments are designed to evaluate and compare the knowledgeability of the LLM across multiple datasets. We illustrate our process on five datasets: MVPKG (covering U.S. legislative, election, diplomatic data, etc.), T-Rex (containing large-scale high-quality alignments between DBpedia abstracts and Wikidata triples), PharmKG8K (biomedical knowledge graph), WD50K (derived from Wikidata statements), and CoDEx-S (extracted from Wikidata and Wikipedia). Table 4 provides the dataset statistics.
  • MVPKG (Mou et al., 2024): The MVPKG dataset encompasses U.S. legislative, election, and diplomatic data as well as conceptual knowledge from Wikidata. It originally contains 1,857,410 triplets, 137,117 entities, and 602 relations. Due to scale considerations, we extract the largest strongly connected component, which comprises 255,697 triplets, 9,055 entities, and 602 relations. The MVPKG dataset had a temporal attribute and was evaluated with the temporal component included and excluded. For each triplet, two prompts are generated (with time and without time). Consequently, each entity in MVPKG is assigned two knowledgeability scores corresponding to the two prompt variants for further analysis of the effect of inclusion of temporal information. All other datasets have only one knowledgeability score due to lack of temporal attributes.
  • T-Rex (Elsahar et al., 2018): The T-Rex dataset is constructed from Wikipedia abstracts aligned with Wikidata entities in English. It contains 6,566,790 unique triplets; the largest connected component comprises 193,781 triplets, 46,891 entities, and 423 relations.
  • PharmKG8K (Zheng et al., 2021): The PharmKG8K multi-relational, attributed biomedical KG, composed of around 500,000 individual interconnections between genes, drugs, and diseases, with 29 relation types over a vocabulary of around 8000 disambiguated entities. Given the scope of the dataset, we used a strongly connected component of 98,537 edges, 6,877 entities, and 29 relations.
  • WD50k (Galkin et al., 2020): The WD50K dataset was created using the Wikidata RDF dump of August 2019. It has 233,838 edges and 41,334 entities. Since being extracted from Wikidata, there were 14,858 triplets common between the WD50K dataset and the T-Rex largest connected component selected. These were removed to make sure that common triplets were not overshadowing the result comparison between these datasets. Following that, the largest strongly connected component was selected for experimental purposes. This LCC had 34,208 edges, 5,140 entities, and 193 relations.
  • CoDEx-S (Safavi and Koutra, 2020): CoDEx is a collection of knowledge graph completion datasets extracted from Wikidata/Wikipedia, comprising three subsets of varying sizes. We select CoDEx-S due to its high proportion of triplets involving the “occupation” relation, which poses greater challenges for LLMs, since individuals may hold multiple occupations. CoDEx-S contains 36,543 triplets, 2,034 entities, and 42 relations.